# Reanimating Images using Neural Representations of Dynamic Stimuli

### v2

@@ 4.3 Identifying Brain Regions Tuned to Dynamic Features
Figure 7: Brain Encoding Prediction Performance. We analyze the ability of visual encoders (VideoMAE [1] and CLIP ConvNeXt) to predict (encode) fMRI brain activity from visual stimuli. Overall we find that self-supervised video models, namely VideoMAE and VideoMAE Large, are best at predicting fMRI brain activity. We present details on which brain regions are best predicted by VideoMAE Large in Figure 8. For static image models, semantically-supervised models, namely CLIP ConvNeXt, perform best (consistent with [8]). Finally, we find that VC-1 performs best among embodied AI models. We compare the relative encoding performance of the best video (VideoMAE) and static image (CLIP ConvNeXt) models in Figure 9. Error bars represent the standard error of the mean.
@@
Within image encoding models, we find that CLIP ConvNeXt predicts fMRI brain activity best across all participants (Pearson $r$’s for individual participants, S1: $0.219$, S2: $0.263$, S3: $0.272$, all $p\ll 0.001$), corroborating prior results [8] that guidance from semantics is useful for visual brain prediction. Incorporating temporal information by averaging embeddings over multiple frames from encoding models that are not trained to be aligned over time does not improve performance.
@@
Figure 8: VideoMAE voxel-wise prediction performance. We plot the voxel-wise fMRI voxel encoding accuracy of the best encoding model, VideoMAE Large, quantified as the Pearson correlation between measured and predicted responses. Please refer to the supplemental for a labeled fMRI data parcellation of brain regions. Note that the majority of the voxels in visual cortex are predicted well by VideoMAE. Interestingly, we find that VideoMAE predicts areas TPOJ1, STV, PGi, and 7PC (all subjects, Left Hemisphere, Lateral View) particularly well. These brain areas process visual and motor information and are important for high-level brain functions, including motor planning and theory of mind [89]. This suggests that VideoMAE is capturing dynamic representations relevant for motor planning.
Figure 9: Voxel-wise prediction difference between VideoMAE and CLIP ConvNeXt. We plot the voxel-wise difference of fMRI voxel encoding accuracy between the best encoding model, VideoMAE Large, and the best performing static image encoding model, CLIP ConvNeXt, quantified as the Pearson correlation between measured and predicted responses. fMRI voxels where VideoMAE predicts better are in blue and where CLIP ConvNeXt predicts better are in red. Note that the majority of visual cortex is better predicted by VideoMAE. We also find that VideoMAE predicts areas 5m, 5mv, 23c, 4, 24dd, and 24dv much better than CLIP ConvNeXt (S1 and S2, Left and Right Hemispheres, Medial View). These areas are part of the somatosensory cortex and integrate visuomotor and somatosensory (tactile, proprioceptive information) information for precise motor control and spatial awareness [90, 91]. This suggests that VideoMAE is capturing action representations which are less prevalent in static image models.
@@ 5 Conclusion
Limitations and Future Work. Our methods successfully decode motion of salient objects from fMRI brain activity on the Dynamic Vision Dataset [87]. However, the generalizability of our conclusions to different dataset distributions remains unclear. Future datasets with fine-grained motion can help determine the limits of motion decodability from fMRI brain activity. Furthermore, our subject-specific training approach could be improved by exploring cross-subject alignment methods, as has been shown to improve decoding performance for static stimuli [92, 93].
Conclusion and Impact. To summarize, in this paper we propose to study neural representations of motion by disentangling the modeling of static image features and motion. We show that brain activity can be used to predict visual motion, demonstrate that our predicted motion can animate static images to reconstruct the observed dynamic visual stimuli, and identify regions in the brain that are tuned for dynamic stimuli. Our results lay the groundwork for better understanding how the brain represents visual motion, and may help researchers construct better brain-machine interfaces.
Supplementary Material:
Reanimating Images using Neural Representations of Dynamic Stimuli
@@ A.3 Decoding Models: Visualizations
Our website https://sites.google.com/view/neural-dynamics/home has multiple reanimation examples as videos. For each example, we show the ground truth video, the reanimated video using the flow predicted from the ground truth initial frame, and the reanimated video using the flow predicted from the initial frame generated by MindVideo [3]. The MindVideo initial frame represents the diffusion image as predicted by MindVideo from fMRI data. As such, this frame is often quite different in content from the ground truth, although the objects in each are in similar locations. However, the motion predicted by our model is still consistent with ground truth motion, for example, in the first video the jellyfish retracts backwards at the end clip. The same backwards motion can be seen in the reanimated videos for with both the ground truth initial frame and the MindVideo generated initial frame (which happens to be a boat). This establishes that the motion visualized using DragNUWA [4] is not derived solely from the diffusion model, but rather, incorporates our predicted motion.
@@
Figure S2: Static image animation results. Here we see an example of deer running to the right in the ground truth video (first row, “GT”). We show an example of animating the initial frame of the ground truth video by combining the brain conditioned motion prediction with DragNUWA [4] (second row, “Ours + GT”). We show an example of animating the initial frame obtained from fMRI data using MindVideo by combining the brain conditioned motion prediction with DragNUWA (third row, “Ours + MindVideo”) and observe that the creatures run to the right.
Figure S3: Static image animation results. Here we see an example of a soldier running to the left in the ground truth video (first row, “GT”). We show an example of animating the initial frame of the ground truth video by combining the brain conditioned motion prediction with DragNUWA [4] (second row, “Ours + GT”). We show an example of animating the initial frame obtained from fMRI data using MindVideo by combining the brain conditioned motion prediction with DragNUWA (third row, “Ours + MindVideo”) and observe that the child pedals its feet in a similar motion to that of the soldier running.

### v3

@@ 4.3 Identifying Brain Regions Tuned to Dynamic Features
Figure 7: Brain Encoding Prediction Performance. We analyze the ability of visual encoders (VideoMAE [1] and CLIP ConvNeXt) to predict (encode) fMRI brain activity from visual stimuli. Self-supervised video models, VideoMAE and VideoMAE Large, best predict fMRI brain activity. Details on which brain regions are best predicted by VideoMAE Large are shown in Figure 8. For static image models, semantically-supervised models such as CLIP ConvNeXt, perform best (consistent with [7]), while VC-1 performs best among embodied AI models. Figure 9 compares the relative encoding performance of the best video (VideoMAE) and static image (CLIP ConvNeXt) models. Error bars represent the standard error of the mean.
@@
Within image encoding models, CLIP ConvNeXt predicts fMRI brain activity best across all participants (Pearson $r$’s for individual participants, S1: $0.219$, S2: $0.263$, S3: $0.272$, all $p\ll 0.001$), consistent with prior results [7] that alignment with language improves visual brain prediction. Averaging embeddings across frames from temporally unaligned encoding models does not enhance performance.
@@
Figure 8: VideoMAE voxel-wise prediction performance. Voxel-wise fMRI voxel encoding accuracy of the best encoding model, VideoMAE Large, quantified as the Pearson correlation between measured and predicted responses. Refer to the supplemental for a labeled fMRI data parcellation of brain regions. The majority of the voxels in visual cortex are predicted well by VideoMAE. VideoMAE predicts areas TPOJ1, STV, PGi, and 7PC (all subjects, Left Hemisphere, Lateral View) particularly well – brain areas are implicated in processing visual and motor information and appear important for high-level brain functions, including motor planning and theory of mind [88]. Thus, VideoMAE may be capturing dynamic information relevant for motor planning.
Figure 9: Voxel-wise prediction difference between VideoMAE and CLIP ConvNeXt. Voxel-wise difference of fMRI voxel encoding accuracy between the best encoding model, VideoMAE Large, and the best performing static image encoding model, CLIP ConvNeXt, quantified as the Pearson correlation between measured and predicted responses. fMRI voxels where VideoMAE predicts better are in blue and where CLIP ConvNeXt predicts better are in red. The majority of visual cortex is better predicted by VideoMAE. VideoMAE also predicts areas 5m, 5mv, 23c, 4, 24dd, and 24dv much better than CLIP ConvNeXt (S1 and S2, Left and Right Hemispheres, Medial View) – all part of the somatosensory cortex which integrates visuomotor and somatosensory information for precise motor control and spatial awareness [89, 90]. Thus, VideoMAE may be capturing action representations which are less prevalent in static image models.
@@ 5 Conclusion
Limitations and Future Work. BrainNRDS successfully decodes motion of salient objects from fMRI brain activity on the Dynamic Vision Dataset [86]. However, the generalizability of our conclusions to different dataset distributions remains unclear. Future datasets with fine-grained motion can help determine the limits of motion decodability from fMRI brain activity. Furthermore, our subject-specific training approach could be improved by exploring cross-subject alignment methods, as has been shown to improve decoding performance for static stimuli [91, 92].
Conclusion and Impact. To summarize, in this paper we propose to study neural representations of motion by disentangling the modeling of static image features and motion using BrainNRDS. We show that brain activity can be used to predict visual motion, demonstrate that our predicted motion can animate static images to reconstruct the observed dynamic visual stimuli, and identify regions in the brain that are tuned for dynamic stimuli. Our results lay the groundwork for better understanding how the brain represents visual motion.
@@ A.3 Decoding Models: Visualizations
Our website https://brain-nrds.github.io/ has multiple reanimation examples as videos. For each example, we show the ground truth video, the reanimated video using the flow predicted from the ground truth initial frame, and the reanimated video using the flow predicted from the initial frame generated by MindVideo [3]. The MindVideo initial frame represents the diffusion image as predicted by MindVideo from fMRI data. As such, this frame is often quite different in content from the ground truth, although the objects in each are in similar locations. However, the motion predicted by our model is still consistent with ground truth motion, for example, in the first video the jellyfish retracts backwards at the end clip. The same backwards motion can be seen in the reanimated videos for with both the ground truth initial frame and the MindVideo generated initial frame (which happens to be a boat). This establishes that the motion visualized using DragNUWA [4] is not derived solely from the diffusion model, but rather, incorporates our predicted motion.
@@
Figure S2: Static image animation results. Here we see an example of deer running to the right in the ground truth video (first row, “GT"). We show an example of animating the initial frame of the ground truth video by combining the brain conditioned motion prediction with DragNUWA [4] (second row, “Ours + GT"). We show an example of animating the initial frame obtained from fMRI data using MindVideo by combining the brain conditioned motion prediction with DragNUWA (third row, “Ours + MindVideo") and observe that the creatures run to the right.
Figure S3: Static image animation results. Here we see an example of a soldier running to the left in the ground truth video (first row, “GT"). We show an example of animating the initial frame of the ground truth video by combining the brain conditioned motion prediction with DragNUWA [4] (second row, “Ours + GT"). We show an example of animating the initial frame obtained from fMRI data using MindVideo by combining the brain conditioned motion prediction with DragNUWA (third row, “Ours + MindVideo") and observe that the child pedals its feet in a similar motion to that of the soldier running.